OpenStudy (anonymous):

what is the measure of an angle whose supplement is four times its complement?

OpenStudy (campbell_st):

so this is simultaneous equations complementary angles x + y = 90 supplementary angles 4x + y = 180 this will give the the angles....

OpenStudy (anonymous):

so how do you do it ?

OpenStudy (campbell_st):

let y = 90 - x and substitute it into 4x + y = 180 4x + (90 -x) = 180 solve the equation

OpenStudy (anonymous):

its 60 though, the answer

OpenStudy (campbell_st):

well 1 angle is 60... and thats the angle that is the same in both cases... what is the complement of 60 what is the supplement of 60... is the supplement 4x the complement... as the question asks...

OpenStudy (anonymous):

huh?